\begin{thebibliography}{70}
%\cite{Maldacena:1997re}

\bibitem{Arsene:2004fa}
  I.~Arsene {\it et al.}  [BRAHMS Collaboration],
  %``Quark gluon plasma and color glass condensate at RHIC? The perspective
  %from the BRAHMS experiment,''
  Nucl.\ Phys.\  A {\bf 757}, 1 (2005)
  [arXiv:nucl-ex/0410020].
  %%CITATION = NUPHA,A757,1;%%
  
%\cite{Adcox:2004mh}
\bibitem{Adcox:2004mh}
  K.~Adcox {\it et al.}  [PHENIX Collaboration],
  %``Formation of dense partonic matter in relativistic nucleus nucleus
  %collisions at RHIC: Experimental evaluation by the PHENIX  collaboration,''
  Nucl.\ Phys.\  A {\bf 757}, 184 (2005)
  [arXiv:nucl-ex/0410003].
  %%CITATION = NUPHA,A757,184;%%

%\cite{Back:2004je}
\bibitem{Back:2004je}
  B.~B.~Back {\it et al.} [PHOBOS Collaboration],
  %``The PHOBOS perspective on discoveries at RHIC,''
  Nucl.\ Phys.\  A {\bf 757}, 28 (2005)
  [arXiv:nucl-ex/0410022].
  %%CITATION = NUPHA,A757,28;%%
  
  %\cite{Adams:2005dq}
\bibitem{Adams:2005dq}
  J.~Adams {\it et al.}  [STAR Collaboration],
  %``Experimental and theoretical challenges in the search for the quark  gluon
  %plasma: The STAR collaboration's critical assessment of the  evidence from
  %RHIC collisions,''
  Nucl.\ Phys.\  A {\bf 757}, 102 (2005)
  [arXiv:nucl-ex/0501009].
  %%CITATION = NUPHA,A757,102;%%

%\cite{Bjorken:1982tu}
\bibitem{Bjorken:1982tu}
  J.~D.~Bjorken,
%  ``Energy Loss Of Energetic Partons In Quark - Gluon Plasma: Possible
%  Extinction Of High P(T) Jets In Hadron - Hadron Collisions,''
 FERMILAB-PUB-82-059-THY, 1982.
 %CITATION = FERMILAB-PUB-82-059-THY%

%\cite{Gyulassy:1993hr}
\bibitem{Gyulassy:1993hr}
  M.~Gyulassy and X.~n.~Wang,
  %``Multiple collisions and induced gluon Bremsstrahlung in QCD,''
  Nucl.\ Phys.\  B {\bf 420}, 583 (1994)
  [arXiv:nucl-th/9306003].
  %%CITATION = NUPHA,B420,583;%%

%\cite{Baier:1996kr}
\bibitem{Baier:1996kr}
  R.~Baier, Y.~L.~Dokshitzer, A.~H.~Mueller, S.~Peigne and D.~Schiff,
  %``Radiative energy loss of high energy quarks and gluons in a  finite-volume
  %quark-gluon plasma,''
  Nucl.\ Phys.\  B {\bf 483}, 291 (1997)
  [arXiv:hep-ph/9607355].
  %%CITATION = NUPHA,B483,291;%%

%\cite{Baier:2000mf}
\bibitem{Baier:2000mf}
  R.~Baier, D.~Schiff and B.~G.~Zakharov,
  %``Energy loss in perturbative QCD,''
  Ann.\ Rev.\ Nucl.\ Part.\ Sci.\  {\bf 50}, 37 (2000)
  [arXiv:hep-ph/0002198].
  %%CITATION = ARNUA,50,37;%%

%\cite{Gyulassy:2003mc}
\bibitem{Gyulassy:2003mc}
  M.~Gyulassy, I.~Vitev, X.~N.~Wang and B.~W.~Zhang,
  %``Jet quenching and radiative energy loss in dense nuclear matter,''
  arXiv:nucl-th/0302077.
  %%CITATION = NUCL-TH/0302077;%%



%\cite{Dokshitzer:2001zm}
\bibitem{Dokshitzer:2001zm}
  Y.~L.~Dokshitzer and D.~E.~Kharzeev,
  %``Heavy quark colorimetry of QCD matter,''
  Phys.\ Lett.\  B {\bf 519}, 199 (2001)
  [arXiv:hep-ph/0106202].
  %%CITATION = PHLTA,B519,199;%%

%\cite{Armesto:2005iq}
\bibitem{Armesto:2005iq}
  N.~Armesto, A.~Dainese, C.~A.~Salgado and U.~A.~Wiedemann,
  %``Testing the color charge and mass dependence of parton energy loss with
  %heavy-to-light ratios at RHIC and LHC,''
  Phys.\ Rev.\  D {\bf 71}, 054027 (2005)
  [arXiv:hep-ph/0501225].
  %%CITATION = PHRVA,D71,054027;%%

%\cite{Djordjevic:2004nq}
\bibitem{Djordjevic:2004nq}
  M.~Djordjevic, M.~Gyulassy and S.~Wicks,
  %``The charm and beauty of RHIC and LHC,''
  Phys.\ Rev.\ Lett.\  {\bf 94}, 112301 (2005)
  [arXiv:hep-ph/0410372].
  %%CITATION = PRLTA,94,112301;%%

%\cite{Adare:2006nq}
\bibitem{Adare:2006nq}
  A.~Adare {\it et al.}  [PHENIX Collaboration],
  %``Energy Loss and Flow of Heavy Quarks in Au+Au Collisions at sqrt(s_NN) =
  %200 GeV,''
  Phys.\ Rev.\ Lett.\  {\bf 98}, 172301 (2007)
  [arXiv:nucl-ex/0611018].
  %%CITATION = PRLTA,98,172301;%%

%\cite{Abelev:2006db}
\bibitem{Abelev:2006db}
  B.~I.~Abelev {\it et al.}  [STAR Collaboration],
  %``Transverse momentum and centrality dependence of high-pt non-photonic
  %electron suppression in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV,''
  Phys.\ Rev.\ Lett.\  {\bf 98}, 192301 (2007)
  [arXiv:nucl-ex/0607012].
  %%CITATION = PRLTA,98,192301;%%

%\cite{:2008hj}
\bibitem{:2008hj}
  B. I. Abelev  {\it et al.}  [STAR Collaboration],
  %``Charmed hadron production at low transverse momentum in Au+Au collisions at
  %RHIC,''
  arXiv:0805.0364 [nucl-ex].
  %%CITATION = ARXIV:0805.0364;%%

%\cite{Awes:2008qi}
\bibitem{Awes:2008qi}
  T.~C.~Awes  [PHENIX Collaboration],
  %``Highlights from PHENIX - II,''
  arXiv:0805.1636 [nucl-ex].
  %%CITATION = ARXIV:0805.1636;%%

\bibitem{Maldacena:1997re}
  J.~M.~Maldacena,
  %``The large N limit of superconformal field theories and supergravity,''
  Adv.\ Theor.\ Math.\ Phys.\  {\bf 2}, 231 (1998)
  [Int.\ J.\ Theor.\ Phys.\  {\bf 38}, 1113 (1999)]
  [arXiv:hep-th/9711200].
  %%CITATION = IJTPB,38,1113;%%

%\cite{Gubser:1998bc}
\bibitem{Gubser:1998bc}
  S.~S.~Gubser, I.~R.~Klebanov and A.~M.~Polyakov,
  %``Gauge theory correlators from non-critical string theory,''
  Phys.\ Lett.\  B {\bf 428}, 105 (1998)
  [arXiv:hep-th/9802109].
  %%CITATION = PHLTA,B428,105;%%

%\cite{Witten:1998qj}
\bibitem{Witten:1998qj}
  E.~Witten,
  %``Anti-de Sitter space and holography,''
  Adv.\ Theor.\ Math.\ Phys.\  {\bf 2}, 253 (1998)
  [arXiv:hep-th/9802150].
  %%CITATION = 00203,2,253;%%

%\cite{Aharony:1999ti}
\bibitem{Aharony:1999ti}
  O.~Aharony, S.~S.~Gubser, J.~M.~Maldacena, H.~Ooguri and Y.~Oz,
  %``Large N field theories, string theory and gravity,''
  Phys.\ Rept.\  {\bf 323}, 183 (2000)
  [arXiv:hep-th/9905111].
  %%CITATION = PRPLC,323,183;%%

%\cite{Kharzeev:2007wb}
\bibitem{Kharzeev:2007wb}
  D.~Kharzeev and K.~Tuchin,
  %``Bulk viscosity of QCD matter near the critical temperature,''
  arXiv:0705.4280 [hep-ph].
  %%CITATION = ARXIV:0705.4280;%%

%\cite{Karsch:2007jc}
\bibitem{Karsch:2007jc}
  F.~Karsch, D.~Kharzeev and K.~Tuchin,
  %``Universal properties of bulk viscosity near the QCD phase transition,''
  Phys.\ Lett.\  B {\bf 663}, 217 (2008)
  [arXiv:0711.0914 [hep-ph]].
  %%CITATION = PHLTA,B663,217;%%

%\cite{Meyer:2007dy}
\bibitem{Meyer:2007dy}
  H.~B.~Meyer,
  %``A calculation of the bulk viscosity in SU(3) gluodynamics,''
  Phys.\ Rev.\ Lett.\  {\bf 100}, 162001 (2008)
  [arXiv:0710.3717 [hep-lat]].
  %%CITATION = PRLTA,100,162001;%%

%\cite{Benincasa:2005iv}
\bibitem{Benincasa:2005iv}
  P.~Benincasa, A.~Buchel and A.~O.~Starinets,
  %``Sound waves in strongly coupled non-conformal gauge theory plasma,''
  Nucl.\ Phys.\  B {\bf 733}, 160 (2006)
  [arXiv:hep-th/0507026].
  %%CITATION = NUPHA,B733,160;%%

%\cite{Buchel:2007mf}
\bibitem{Buchel:2007mf}
  A.~Buchel,
  %``Bulk viscosity of gauge theory plasma at strong coupling,''
  Phys.\ Lett.\  B {\bf 663}, 286 (2008)
  [arXiv:0708.3459 [hep-th]].
  %%CITATION = PHLTA,B663,286;%%

%\cite{Gubser:2008yx}
\bibitem{Gubser:2008yx}
  S.~S.~Gubser, A.~Nellore, S.~S.~Pufu and F.~D.~Rocha,
  %``Thermodynamics and bulk viscosity of approximate black hole duals to finite
  %temperature quantum chromodynamics,''
  arXiv:0804.1950 [hep-th].
  %%CITATION = ARXIV:0804.1950;%%

%\cite{Policastro:2001yc}
\bibitem{Policastro:2001yc}
  G.~Policastro, D.~T.~Son and A.~O.~Starinets,
  %``The shear viscosity of strongly coupled N = 4 supersymmetric Yang-Mills
  %plasma,''
  Phys.\ Rev.\ Lett.\  {\bf 87}, 081601 (2001)
  [arXiv:hep-th/0104066].
  %%CITATION = PRLTA,87,081601;%%

%\cite{Kovtun:2004de}
\bibitem{Kovtun:2004de}
  P.~Kovtun, D.~T.~Son and A.~O.~Starinets,
  %``Viscosity in strongly interacting quantum field theories from black hole
  %physics,''
  Phys.\ Rev.\ Lett.\  {\bf 94}, 111601 (2005)
  [arXiv:hep-th/0405231].
  %%CITATION = PRLTA,94,111601;%%

%\cite{Herzog:2006gh}
\bibitem{Herzog:2006gh}
  C.~P.~Herzog, A.~Karch, P.~Kovtun, C.~Kozcaz and L.~G.~Yaffe,
  %``Energy loss of a heavy quark moving through N = 4 supersymmetric
  %Yang-Mills plasma,''
  JHEP {\bf 0607}, 013 (2006)
  [arXiv:hep-th/0605158].
  %%CITATION = JHEPA,0607,013;%%

%\cite{Gubser:2006bz}
\bibitem{Gubser:2006bz}
  S.~S.~Gubser,
  %``Drag force in AdS/CFT,''
  Phys.\ Rev.\  D {\bf 74}, 126005 (2006)
  [arXiv:hep-th/0605182].
  %%CITATION = PHRVA,D74,126005;%%
  
%\cite{CasalderreySolana:2006rq}
\bibitem{CasalderreySolana:2006rq}
  J.~Casalderrey-Solana and D.~Teaney,
  %``Heavy quark diffusion in strongly coupled N = 4 Yang Mills,''
  Phys.\ Rev.\  D {\bf 74}, 085012 (2006)
  [arXiv:hep-ph/0605199].
  %%CITATION = PHRVA,D74,085012;%%

%\cite{Horowitz:2008ig}
\bibitem{Horowitz:2008ig}
  W.~A.~Horowitz and M.~Gyulassy,
  %``Testing AdS/CFT Drag and pQCD Heavy Quark Energy Loss,''
  arXiv:0804.4330 [hep-ph].
  %%CITATION = ARXIV:0804.4330;%%
  
%\cite{Liu:2006ug}
\bibitem{Liu:2006ug}
  H.~Liu, K.~Rajagopal and U.~A.~Wiedemann,
  %``Calculating the jet quenching parameter from AdS/CFT,''
  Phys.\ Rev.\ Lett.\  {\bf 97}, 182301 (2006)
  [arXiv:hep-ph/0605178].
  %%CITATION = PRLTA,97,182301;%%

%\cite{Chernicoff:2006hi}
\bibitem{Chernicoff:2006hi}
  M.~Chernicoff, J.~A.~Garcia and A.~Guijosa,
  %``The energy of a moving quark-antiquark pair in an N = 4 SYM plasma,''
  JHEP {\bf 0609}, 068 (2006)
  [arXiv:hep-th/0607089].
  %%CITATION = JHEPA,0609,068;%%

%\cite{Argyres:2006yz}
\bibitem{Argyres:2006yz}
  P.~C.~Argyres, M.~Edalati and J.~F.~Vazquez-Poritz,
  %``Spacelike strings and jet quenching from a Wilson loop,''
  JHEP {\bf 0704}, 049 (2007)
  [arXiv:hep-th/0612157].
  %%CITATION = JHEPA,0704,049;%%

%\cite{Dominguez:2008vd}
\bibitem{Dominguez:2008vd}
  F.~Dominguez, C.~Marquet, A.~H.~Mueller, B.~Wu and B.~W.~Xiao,
  %``Comparing energy loss and $p_{\perp}$-broadening in perturbative QCD with
  %strong coupling $\mathcal{N}=4$ SYM theory,''
  arXiv:0803.3234 [nucl-th].
  %%CITATION = ARXIV:0803.3234;%%

%\cite{Hatta:2008tx}
\bibitem{Hatta:2008tx}
  Y.~Hatta, E.~Iancu and A.~H.~Mueller,
  %``Jet evolution in the N=4 SYM plasma at strong coupling,''
  JHEP {\bf 0805}, 037 (2008)
  [arXiv:0803.2481 [hep-th]].
  %%CITATION = JHEPA,0805,037;%%

%\cite{Mueller:2008zt}
\bibitem{Mueller:2008zt}
  A.~H.~Mueller,
  %``Separating hard and soft scales in hard processes in a QCD plasma,''
  arXiv:0805.3140 [hep-ph].
  %%CITATION = ARXIV:0805.3140;%%

%\cite{Mikhailov:2003er}
\bibitem{Mikhailov:2003er}
  A.~Mikhailov,
  %``Nonlinear waves in AdS/CFT correspondence,''
  arXiv:hep-th/0305196.
  %%CITATION = HEP-TH/0305196;%%

\bibitem{lienard}
A. Li\'enard, L'\'Eclairage \'Electrique {\bf 16}, 5 (1898).

\bibitem{Landau}
L. D. Landau and E. M. Lifshitz, "The classical theory of fields", Pergamon Press, 1975.

\bibitem{Jackson}
J. D. Jackson, "Classical electrodynamics", John Wiley $\&$ Sons, Inc., 1975.  

 
 %\cite{Dokshitzer:1977sg}
\bibitem{Dokshitzer:1977sg}
  Y.~L.~Dokshitzer,
  %``Calculation Of The Structure Functions For Deep Inelastic Scattering And E+
  %E- Annihilation By Perturbation Theory In Quantum Chromodynamics. (In
  %Russian),''
  Sov.\ Phys.\ JETP {\bf 46}, 641 (1977)
  [Zh.\ Eksp.\ Teor.\ Fiz.\  {\bf 73}, 1216 (1977)].
  %%CITATION = ZETFA,73,1216;%%
%\cite{Altarelli:1977zs}

\bibitem{Altarelli:1977zs}
  G.~Altarelli and G.~Parisi,
  %``Asymptotic Freedom In Parton Language,''
  Nucl.\ Phys.\  B {\bf 126}, 298 (1977).
  %%CITATION = NUPHA,B126,298;%% 

%\cite{Gribov:1972ri}
\bibitem{Gribov:1972ri}
  V.~N.~Gribov and L.~N.~Lipatov,
  %``Deep Inelastic E P Scattering In Perturbation Theory,''
  Sov.\ J.\ Nucl.\ Phys.\  {\bf 15}, 438 (1972)
  [Yad.\ Fiz.\  {\bf 15}, 781 (1972)].
  %%CITATION = YAFIA,15,781;%%

%\cite{Mandal:2002fs}
\bibitem{Mandal:2002fs}
  G.~Mandal, N.~V.~Suryanarayana and S.~R.~Wadia,
  %``Aspects of semiclassical strings in AdS(5),''
  Phys.\ Lett.\  B {\bf 543}, 81 (2002)
  [arXiv:hep-th/0206103].
  %%CITATION = PHLTA,B543,81;%%

%\cite{Bena:2003wd}
\bibitem{Bena:2003wd}
  I.~Bena, J.~Polchinski and R.~Roiban,
  %``Hidden symmetries of the AdS(5) x S**5 superstring,''
  Phys.\ Rev.\  D {\bf 69}, 046002 (2004)
  [arXiv:hep-th/0305116].
  %%CITATION = PHRVA,D69,046002;%%

%\cite{Sin:2004yx}
\bibitem{Sin:2004yx}
  S.~J.~Sin and I.~Zahed,
  %``Holography of radiation and jet quenching,''
  Phys.\ Lett.\  B {\bf 608}, 265 (2005)
  [arXiv:hep-th/0407215].
  %%CITATION = PHLTA,B608,265;%%

%\cite{Chernicoff:2008sa}
\bibitem{Chernicoff:2008sa}
  M.~Chernicoff and A.~Guijosa,
  %``Acceleration, Energy Loss and Screening in Strongly-Coupled Gauge
  %Theories,''
  arXiv:0803.3070 [hep-th].
  %%CITATION = ARXIV:0803.3070;%%

%\cite{'t Hooft:1981ht}
\bibitem{'t Hooft:1981ht}
  G.~'t Hooft,
  %``Topology Of The Gauge Condition And New Confinement Phases In Nonabelian
  %Gauge Theories,''
  Nucl.\ Phys.\  B {\bf 190}, 455 (1981).
  %%CITATION = NUPHA,B190,455;%%

%\cite{Korthals Altes:2006gx}
\bibitem{Korthals Altes:2006gx}
  C.~P.~Korthals Altes,
  %``Magnetic monopoles in hot QCD,''
  arXiv:hep-ph/0607154.
  %%CITATION = HEP-PH/0607154;%%

%\cite{Shuryak:2007qs}
\bibitem{Shuryak:2007qs}
  E.~Shuryak,
  %``Emerging theory of strongly coupled quark-gluon plasma,''
  arXiv:hep-ph/0703208.
  %%CITATION = HEP-PH/0703208;%%

\bibitem{born}
M. Born and L. Infeld, Nature {\bf 132}, 1004 (1933).

%\cite{Gribov:1999ui}
\bibitem{Gribov:1999ui}
  V.~N.~Gribov,
  %``The theory of quark confinement,''
  Eur.\ Phys.\ J.\  C {\bf 10}, 91 (1999)
  [arXiv:hep-ph/9902279].
  %%CITATION = EPHJA,C10,91;%%

%\cite{Dokshitzer:2004ie}
\bibitem{Dokshitzer:2004ie}
  Y.~L.~Dokshitzer and D.~E.~Kharzeev,
  %``The Gribov conception of quantum chromodynamics,''
  Ann.\ Rev.\ Nucl.\ Part.\ Sci.\  {\bf 54}, 487 (2004)
  [arXiv:hep-ph/0404216].
  %%CITATION = ARNUA,54,487;%%

%\cite{Kharzeev:2005iz}
\bibitem{Kharzeev:2005iz}
  D.~Kharzeev and K.~Tuchin,
  %``From color glass condensate to quark gluon plasma through the event
  %horizon,''
  Nucl.\ Phys.\  A {\bf 753}, 316 (2005)
  [arXiv:hep-ph/0501234].
  %%CITATION = NUPHA,A753,316;%%

%\cite{Kharzeev:2006zm}
\bibitem{Kharzeev:2006zm}
  D.~Kharzeev, E.~Levin and K.~Tuchin,
  %``Multi-particle production and thermalization in high-energy QCD,''
  Phys.\ Rev.\  C {\bf 75}, 044903 (2007)
  [arXiv:hep-ph/0602063].
  %%CITATION = PHRVA,C75,044903;%%

%\cite{Lappi:2006fp}
\bibitem{Lappi:2006fp}
  T.~Lappi and L.~McLerran,
  %``Some features of the glasma,''
  Nucl.\ Phys.\  A {\bf 772}, 200 (2006)
  [arXiv:hep-ph/0602189].
  %%CITATION = NUPHA,A772,200;%%

%\cite{Fries:2006pv}
\bibitem{Fries:2006pv}
  R.~J.~Fries, J.~I.~Kapusta and Y.~Li,
  %``Near-fields and initial energy density in the color glass condensate
  %model,''
  arXiv:nucl-th/0604054.
  %%CITATION = NUCL-TH/0604054;%%

%\cite{Iwazaki:2007es}
\bibitem{Iwazaki:2007es}
  A.~Iwazaki,
  %``Thermalization of Color Gauge Fields in High Energy Heavy Ion Collisions,''
  Phys.\ Rev.\  C {\bf 77}, 034907 (2008)
  [arXiv:0712.1405 [hep-ph]].
  %%CITATION = PHRVA,C77,034907;%%

%\cite{Fujii:2008dd}
\bibitem{Fujii:2008dd}
  H.~Fujii and K.~Itakura,
  %``Expanding color flux tubes and instabilities,''
  arXiv:0803.0410 [hep-ph].
  %%CITATION = ARXIV:0803.0410;%%

%\cite{Gribov:1984tu}
\bibitem{Gribov:1984tu}
  L.~V.~Gribov, E.~M.~Levin and M.~G.~Ryskin,
  %``Semihard Processes In QCD,''
  Phys.\ Rept.\  {\bf 100}, 1 (1983).
  %%CITATION = PRPLC,100,1;%%

%\cite{Blaizot:1987nc}
\bibitem{Blaizot:1987nc}
  J.~P.~Blaizot and A.~H.~Mueller,
  %``The Early Stage Of Ultrarelativistic Heavy Ion Collisions,''
  Nucl.\ Phys.\  B {\bf 289}, 847 (1987).
  %%CITATION = NUPHA,B289,847;%%

%\cite{McLerran:1993ni}
\bibitem{McLerran:1993ni}
  L.~D.~McLerran and R.~Venugopalan,
  %``Computing quark and gluon distribution functions for very large nuclei,''
  Phys.\ Rev.\  D {\bf 49}, 2233 (1994)
  [arXiv:hep-ph/9309289];
  %%CITATION = PHRVA,D49,2233;%%
%``Green's functions in the color field of a large nucleus,''
  Phys.\ Rev.\  D {\bf 50}, 2225 (1994)
  [arXiv:hep-ph/9402335].
  %%CITATION = PHRVA,D50,2225;%%  
%\cite{Kharzeev:2000ef}
\bibitem{Kharzeev:2000ef}
  D.~E.~Kharzeev, Y.~V.~Kovchegov and E.~Levin,
  %``QCD instantons and the soft pomeron,''
  Nucl.\ Phys.\  A {\bf 690}, 621 (2001)
  [arXiv:hep-ph/0007182].
  %%CITATION = NUPHA,A690,621;%%

%\cite{Shuryak:2001cp}
\bibitem{Shuryak:2001cp}
  E.~V.~Shuryak,
  %``The instanton/sphaleron mechanism of gluon production and high energy
  %heavy ion collisions at RHIC,''
  Phys.\ Lett.\  B {\bf 515}, 359 (2001)
  [arXiv:hep-ph/0101269].
  %%CITATION = PHLTA,B515,359;%%

%\cite{Kharzeev:2001vs}
\bibitem{Kharzeev:2001vs}
  D.~E.~Kharzeev, Y.~V.~Kovchegov and E.~Levin,
  %``Instantons in the saturation environment,''
  Nucl.\ Phys.\  A {\bf 699}, 745 (2002)
  [arXiv:hep-ph/0106248].
  %%CITATION = NUPHA,A699,745;%%

\bibitem{Kharzeev:2001ev}
  D.~Kharzeev, A.~Krasnitz and R.~Venugopalan,
  %``Anomalous chirality fluctuations in the initial stage of heavy ion
  %collisions and parity odd bubbles,''
  Phys.\ Lett.\  B {\bf 545}, 298 (2002)
  [arXiv:hep-ph/0109253].
  %%CITATION = PHLTA,B545,298;%%

%\cite{Janik:2002nk}
\bibitem{Janik:2002nk}
  R.~A.~Janik, E.~Shuryak and I.~Zahed,
  %``Prompt multi-gluon production in high energy collisions from singular
  %Yang-Mills solutions,''
  Phys.\ Rev.\  D {\bf 67}, 014005 (2003)
  [arXiv:hep-ph/0206005].
  %%CITATION = PHRVA,D67,014005;%%

%\cite{Kharzeev:1998kz}
\bibitem{Kharzeev:1998kz}
  D.~Kharzeev, R.~D.~Pisarski and M.~H.~G.~Tytgat,
  %``Possibility of spontaneous parity violation in hot {QCD},''
  Phys.\ Rev.\ Lett.\  {\bf 81}, 512 (1998)
  [arXiv:hep-ph/9804221].
  %%CITATION = PRLTA,81,512;%%

%\cite{Lee:1973iz}
\bibitem{Lee:1973iz}
  T.~D.~Lee,
  %``A Theory of Spontaneous T Violation,''
  Phys.\ Rev.\  D {\bf 8}, 1226 (1973).
  %%CITATION = PHRVA,D8,1226;%%

%\cite{Morley:1983wr}
\bibitem{Morley:1983wr}
  P.~D.~Morley and I.~A.~Schmidt,
  %``Strong P, Cp, T Violations In Heavy Ion Collisions,''
  Z.\ Phys.\  C {\bf 26}, 627 (1985).
  %%CITATION = ZEPYA,C26,627;%%

%\cite{Kharzeev:2004ey}
\bibitem{Kharzeev:2004ey}
  D.~Kharzeev,
  %``Parity violation in hot QCD: Why it can happen, and how to look for it,''
  Phys.\ Lett.\  B {\bf 633}, 260 (2006)
  [arXiv:hep-ph/0406125].
  %%CITATION = PHLTA,B633,260;%%

%\cite{Kharzeev:2007tn}
\bibitem{Kharzeev:2007tn}
  D.~Kharzeev and A.~Zhitnitsky,
  %``Charge separation induced by P-odd bubbles in QCD matter,''
  Nucl.\ Phys.\  A {\bf 797}, 67 (2007)
  [arXiv:0706.1026 [hep-ph]].
  %%CITATION = NUPHA,A797,67;%%

%\cite{Kharzeev:2007jp}
\bibitem{Kharzeev:2007jp}
  D.~E.~Kharzeev, L.~D.~McLerran and H.~J.~Warringa,
  %``The effects of topological charge change in heavy ion collisions: 'Event by
  %event P and CP violation',''
  Nucl.\ Phys.\  A {\bf 803}, 227 (2008)
  [arXiv:0711.0950 [hep-ph]].
  %%CITATION = NUPHA,A803,227;%%

%\cite{Warringa:2008kv}
\bibitem{Warringa:2008kv}
  H.~J.~Warringa,
  %``Implications of CP-violating transitions in hot quark matter on heavy ion
  %collisions,''
  arXiv:0805.1384 [hep-ph].
  %%CITATION = ARXIV:0805.1384;%%

%\cite{Voloshin:2008kc}
\bibitem{Voloshin:2008kc}
  S.~A.~Voloshin,
  %``Anisotropic flow: Achievements, Difficulties, Expectations,''
  arXiv:0805.1351 [nucl-ex].
  %%CITATION = ARXIV:0805.1351;%%

%\cite{Dumitru:2008wn}
\bibitem{Dumitru:2008wn}
  A.~Dumitru, F.~Gelis, L.~McLerran and R.~Venugopalan,
  %``Glasma flux tubes and the near side ridge phenomenon at RHIC,''
  arXiv:0804.3858 [hep-ph].
  %%CITATION = ARXIV:0804.3858;%%

%\cite{Shuryak:2002ai}
\bibitem{Shuryak:2002ai}
  E.~V.~Shuryak and I.~Zahed,
  %``Jet quenching in high energy heavy ion collisions by QCD  synchrotron-like
  %radiation,''
  Phys.\ Rev.\  D {\bf 67}, 054025 (2003)
  [arXiv:hep-ph/0207163].
  %%CITATION = PHRVA,D67,054025;%%

\bibitem{schwinger}
J. Schwinger, Phys. Rev. {\bf 75}, 1912 (1949).

\bibitem{chuk}
I. Ya. Pomeranchuk, J. Phys. {\bf II}, 65 (1940).

%\cite{Kharzeev:2003sk}
\bibitem{Kharzeev:2003sk}
  D.~Kharzeev and K.~Tuchin,
  %``Open charm production in heavy ion collisions and the color glass
  %condensate,''
  Nucl.\ Phys.\  A {\bf 735}, 248 (2004)
  [arXiv:hep-ph/0310358].
  %%CITATION = NUPHA,A735,248;%%

%\cite{Kharzeev:2000ph}
\bibitem{Kharzeev:2000ph}
  D.~Kharzeev and M.~Nardi,
  %``Hadron production in nuclear collisions at RHIC and high density QCD,''
  Phys.\ Lett.\  B {\bf 507}, 121 (2001)
  [arXiv:nucl-th/0012025];
  %%CITATION = PHLTA,B507,121;%%
%\cite{Kharzeev:2001gp}
  D.~Kharzeev and E.~Levin,
  %``Manifestations of high density QCD in the first RHIC data,''
  Phys.\ Lett.\  B {\bf 523}, 79 (2001)
  [arXiv:nucl-th/0108006].
  %%CITATION = PHLTA,B523,79;%%

%\cite{Kharzeev:2004if}
\bibitem{Kharzeev:2004if}
  D.~Kharzeev, E.~Levin and M.~Nardi,
  %``Color glass condensate at the LHC: Hadron multiplicities in p p, p A  and A
  %A collisions,''
  Nucl.\ Phys.\  A {\bf 747}, 609 (2005)
  [arXiv:hep-ph/0408050].
  %%CITATION = NUPHA,A747,609;%%
\end{thebibliography}

\documentclass[12pt,preprintnumbers,aps,amssymb,nofootinbib,amsmath]{revtex4}
\usepackage{epsfig,epsf}
\usepackage{graphicx}
\usepackage{epstopdf}
\newcommand{\beq}{\begin{equation}}
\newcommand{\eeq}{\end{equation}}
\def\eq#1{{(\ref{#1})}}

\newcommand{\as}{\alpha_s}

\newcommand{\ben}{\begin{eqnarray*}}
\newcommand{\een}{\end{eqnarray*}}

\begin{document} 

\preprint{BNL-NT-08/19}

\title{Universal upper bound on the energy of a parton escaping from the strongly coupled quark-gluon matter}

\author{Dmitri E. Kharzeev}
 
\affiliation{Nuclear Theory Group,\\
Department of Physics, Brookhaven National Laboratory,\\
Upton, New York 11973-5000, USA
}

\date{\today}
\pacs{}

\begin{abstract} 
It has been shown through the AdS/CFT correspondence that the energy loss of a fast quark in a strongly coupled ${\cal N}=4$ SUSY Yang--Mills matter in the large $N$ limit is given by the classical Li\'enard formula. I demonstrate that under quite natural assumptions about the dynamics of heavy ion collisions this leads to a universal (i.e. independent of the initial parton energy, but dependent on flavor and  centrality) upper bound on the energy of the partons escaping from the plasma. This bound is a Yang--Mills analog of the Pomeranchuk bound in classical electrodynamics, where it is a consequence of radiation in a strong external field acting on a relativistic charge.   
Since as a result the massive constituent partons are slowed down to a velocity $v < c$, 
 the angular distribution of the emitted radiation exhibits a broad "dead cone". 
If the properties of conformal and QCD matter at strong coupling are qualitatively similar, the existence of this universal upper bound would have dramatic implications for heavy ion experiments. 
\end{abstract}
\maketitle 
One of the most striking results coming from RHIC heavy ion program is the observed suppression of high transverse momentum particles \cite{Arsene:2004fa,Adcox:2004mh,Back:2004je,Adams:2005dq}. Such a suppression resulting from the parton energy loss has been predicted  \cite{Bjorken:1982tu,Gyulassy:1993hr,Baier:1996kr} as a signature of the formation of dense quark-gluon matter. 
The radiative energy loss \cite{Gyulassy:1993hr,Baier:1996kr} has been found to describe the magnitude of the observed suppression, see for example reviews  \cite{Baier:2000mf,Gyulassy:2003mc}. 
A definitive test of this mechanism can be performed with heavy quarks \cite{Dokshitzer:2001zm} -- since heavy quarks move with velocity $v < c$, the induced radiation must be depleted due to the "dead cone" effect -- the vanishing of radiation intensity in the forward direction. This reduces the amount of energy loss, and results in weaker suppression for charm and bottom quarks. 
The heavy-to-light ratios at high transverse momentum have thus been predicted to exceed unity \cite{Dokshitzer:2001zm}; this conclusion survives after the differences in the production mechanisms and fragmentation functions for heavy and light quarks are taken into account \cite{Armesto:2005iq,Djordjevic:2004nq}. 

It thus came as a surprise when RHIC experiments \cite{Abelev:2006db,Adare:2006nq} observed a strong suppression of the high transverse momentum electrons originating from the decays of charmed and beautiful hadrons. The magnitude of the observed suppression indicates that heavy and light quarks are attenuated very similarly in hot QCD matter, in sharp disagreement with the theoretical expectations. This remains true even in the range of transverse momenta that may be dominated by the decays of $b$ quarks \cite{Abelev:2006db,Adare:2006nq}. Moreover, the momenta of the heavy quarks also seem to be strongly deflected by the medium, as is evidenced by the observed elliptical flow (azimuthal anisotropy of the produced heavy quarks with respect to the reaction plane) \cite{Adare:2006nq,:2008hj,Awes:2008qi}.

While it is not yet entirely clear that a perturbative approach cannot be reconciled with the data, the heavy quark energy loss puzzle as well as some other observations in high transverse momenta phenomena \cite{Arsene:2004fa,Adcox:2004mh,Back:2004je,Adams:2005dq} provide an ample  motivation to think about possible non--perturbative effects. Unfortunately, the set of tools that can be used to address the real--time dynamics of QCD in the strong coupling domain is limited. 
Therefore any information about the behavior of gauge theories at strong coupling is very valuable and may provide a hint on how to deal with non-perturbative effects in QCD matter.

\medskip

There has been a major breakthrough in the understanding of strong coupling dynamics of ${\cal N}=4$ SUSY Yang-Mills theory made possible by the AdS/CFT correspondence \cite{Maldacena:1997re,Gubser:1998bc,Witten:1998qj,Aharony:1999ti}. 
In particular, the strong coupling dynamics on the gauge theory side appears dual to the classical supergravity in $AdS_5 \times S_5$ space. 
Of course, QCD and ${\cal N}=4$ SUSY Yang-Mills are very different theories, and conformal invariance of the latter is a crucial 
property which determines the metric of the $AdS_5$ space. Conformal invariance results in the absence of confinement and asymptotic freedom in  ${\cal N}=4$ SUSY Yang-Mills theory. 
Thermodynamical and transport properties of ${\cal N}=4$ SUSY Yang-Mills are also different; in particular the bulk viscosity (related to the scale anomaly of QCD \cite{Kharzeev:2007wb,Karsch:2007jc,Meyer:2007dy}) vanishes unless the conformal symmetry is broken  \cite{Benincasa:2005iv,Buchel:2007mf,Gubser:2008yx}. Nevertheless one hopes that at least some aspects of the plasma behavior at strong coupling may be universal for ${\cal N}=4$ SUSY Yang-Mills and QCD at temperatures higher  than the deconfinement and chiral symmetry restoration temperature (but not much higher so that the coupling is still strong). 
Various properties of the plasma have been considered through the AdS/CFT correspondence, including the computation of  the shear viscosity-to-entropy ratio in the strong coupling limit \cite{Policastro:2001yc,Kovtun:2004de}. The Langevin dynamics of drag force acting on a massive quark traversing the ${\cal N}=4$ SUSY Yang-Mills plasma has been determined by solving the dual problem of string trailing in the ${\rm AdS_5}$ Schwarzschild background \cite{Herzog:2006gh,Gubser:2006bz,CasalderreySolana:2006rq}; a comparison of AdS/CFT drag and pQCD predictions for observables in heavy ion collisions has been recently performed in \cite{Horowitz:2008ig}. A different approach proposed in \cite{Liu:2006ug} aims at matching ${\cal N}=4$ SUSY Yang-Mills calculations to the perturbative QCD. The related problem has been discussed in a number of papers, including \cite{Chernicoff:2006hi,Argyres:2006yz,Dominguez:2008vd,Hatta:2008tx,Mueller:2008zt}.

\medskip
 
In an illuminating paper \cite{Mikhailov:2003er},  it has been shown by Mikhailov that the energy lost by an accelerating external source in ${\cal N}=4$ SUSY Yang-Mills theory in the large $N$ limit is given by the following formula:
\beq\label{lien}
E_{rad}(T) = \frac{\sqrt{\lambda}}{2 \pi} \int_{- \infty}^T \ d t \ \frac{\vec a^2 - (\vec v \times \vec a)^2}{(1 - \vec v^2)^3},
\eeq
where $\lambda=g^2 N$ is 't Hooft coupling, $\vec v$ is 
velocity and $\vec a = \dot{\vec v}$ is acceleration of the charge; we put $c = 1$. Remarkably, with the substitution 
\beq \label{subst}
\frac{\sqrt{\lambda}}{2 \pi} \leftrightarrow \frac{2 e^2}{3}
\eeq
this is exactly the Li\'enard formula \cite{lienard} dating back to 1898 which describes the radiative energy loss in external electromagnetic fields in classical electrodynamics \cite{Landau,Jackson}. In the case of electrodynamics, the linear dependence of the radiated energy on the path length is a consequence of the linearity of Maxwell equations: the emission of electromagnetic radiation at any time is not affected by the previously emitted radiation.  

Yang-Mills equations however are non-linear, and one expects, and indeed finds at weak coupling \cite{Baier:1996kr}, a non-linear dependence of the radiated energy on the path length.  This non-linear dependence is a purely non-Abelian effect: the amplitude of radiation at any given time  is affected by the previously emitted radiation. Such non-linear, non-local dependence on the path traversed by the charge is also responsible for the evolution of jet structure in the vacuum as described by DGLAP equations \cite{Dokshitzer:1977sg,Altarelli:1977zs,Gribov:1972ri}, well established and tested experimentally.  The linear local dependence of the energy lost by the quark in the strong coupling regime as given by \eq{lien} is thus a highly non-trivial result. 

The method of \cite{Mikhailov:2003er} is based on considering a Wilson loop with boundaries given by the external quark and anti-quark sources; in $AdS_5\times S_5$ space, 
the quark and anti-quark are connected by a classical string with two boundaries. The worldsheet of this string is an extremal surface in $AdS_2 \subset AdS_5$, and the energy of the accelerating quark is lost to the excitations on this surface -- non-linear waves. The propagation  of the nonlinear 
wave on the string worldsheet in the large $N$ limit is described by  
the classical sigma model \cite{Mikhailov:2003er}, and the linear dependence of the final result \eq{lien} on the path stems from the integrability of this model \cite{Mandal:2002fs,Bena:2003wd}\footnote{If it appears that this linear Abelian-like formula holds in QCD at strong coupling, the reason may be the conjectured quasi-Abelian dominance proposed by 't Hooft \cite{'t Hooft:1981ht}; this would imply a dynamical role for magnetic monopoles, also in the plasma \cite{Korthals Altes:2006gx,Shuryak:2007qs}.}.  In the large $N$ limit the interactions with closed strings are suppressed, and so the entire lost energy can be attributed to the non--linear wave on the extremal surface stretched between the quark and anti-quark. It is yet unclear (at least to the present author) what excitation corresponds to this wave in the dual Minkowski space gauge theory language; nevertheless below we will attempt to give a qualitative picture based on the analogy with electrodynamics of strong fields.   

\medskip

It should be noted that once the linear, local dependence on the path is established, Lorentz  invariance and dimensional counting completely determine the structure of the formula for the lost energy \cite{Landau,Jackson} -- it has to be proportional to the Li\'enard formula in classical electrodynamics.
Indeed, the relativistic expression for the energy-momentum vector $P^{\mu}$ of the emitted radiation  reads  \cite{Landau,Jackson}, with the substitution \eq{subst}
\beq\label{relat}
P^{\mu} = \frac{\sqrt{\lambda}}{2 \pi} \ \int \frac{d^2 x_\rho}{d \tau^2}\  \frac{d^2 x^\rho}{d \tau^2} \ d x^{\mu},
\eeq
where $\tau$ is the proper time; this is exactly Mikhailov's result \cite{Mikhailov:2003er}. 
 This means that the coincidence of the result \eq{lien} with the Li\'enard formula cannot be considered as an evidence that the mechanism of energy loss at strong coupling is classical radiation; having this in mind, we will nevertheless for simplicity use the familiar language of electrodynamics and refer to the flow of the emitted energy and momentum described by \eq{lien} and \eq{relat} as "radiation". 

The result of Mikhailov \cite{Mikhailov:2003er} has been verified and extended by 
Sin and Zahed \cite{Sin:2004yx} and by Chernicoff and Guijosa \cite{Chernicoff:2008sa}. Sin and Zahed in particular have argued that high momentum partons in a strongly coupled plasma would not be able to penetrate beyond the distance of $1/\pi T$ \cite{Sin:2004yx}. Chernicoff and Guijosa have derived the expression for the dispersion relation of moving quark, and have considered also the case of finite temperature  \cite{Chernicoff:2008sa}.
 
\medskip  
 
 In this paper I will show that under quite natural assumptions about the evolution of gauge fields in heavy ion collisions, the result \eq{lien} implies the existence of a universal (i.e. independent of the initial energy, but dependent on the mass of the parton and on centrality of the collision) upper bound on the final energy of the parton escaping from the strongly coupled matter. 
 
 Let us begin by considering the special case of acceleration parallel to the velocity, $\vec a \parallel \vec v$. Introducing 
 $\gamma = 1/\sqrt{1-v^2}$ and the momentum $\vec p = \gamma m \vec v$,  we get from \eq{lien} or \eq{relat} 
the expression for the power of radiation (radiated energy per unit time):
 \beq
\frac{d E_{rad}}{d t} = \frac{\sqrt{\lambda}}{2 \pi}  \ \frac{1}{\gamma^2 m^2}\ \left(\frac{d \vec p}{d \tau}\right)^2,
\eeq
Since $d \tau = d t / \gamma$, and $d \vec p / d \tau = \gamma\ d \vec p / d t = \gamma \vec F$, where $\vec F$ is the force acting on the charge we get 
 \beq\label{radpar}
\frac{d E_{rad}}{d t} = \frac{\sqrt{\lambda}}{2 \pi} \frac{1}{m^2}\ \ \vec F^2.
\eeq 
The radiation power $d E_{rad}/d t$ in the case of $\vec a \parallel \vec v$ is thus independent of the energy of the charge and is determined only by the magnitude of the external force, as is well-known from
 classical electrodynamics \cite{Landau,Jackson}. The situation when $\vec a \parallel \vec v$ is encountered for example when a quark jet propagates in the vacuum and is slowed down by the force of the string $F = dp/dt = dE/dx = \sigma$, where $\sigma$ is the string tension. In this case the rate of quark energy loss as given by \eq{radpar} does not depend on energy.
 
The expressions found in  \cite{Chernicoff:2008sa} for the energy and momentum of the propagating quark at finite mass diverge as the value of the external force approaches 
\beq\label{critical}
F_{crit} = \frac{2 \pi}{\sqrt{\lambda}}\ m^2.
\eeq
It is natural to interpret this in analogy with electrodynamics as the  critical value of the force capable to produce quark--antiquark pairs from the vacuum. The limiting value of the field strength can be incorporated in a non-linear generalization of electrodynamics uniquely determined by Lorentz invariance and causality proposed by Born and Infeld \cite{born}. Indeed, on the string theory side the limiting value $F_{crit}$ enters the Born-Infeld lagrangian on the D7 brane; once $F \geq F_{crit}$, the creation of open strings becomes energetically favorable, and the system becomes unstable 
\cite{Chernicoff:2008sa}. The mass $m$ in \eq{critical} should be thought of as the constituent quark mass related to the D7-brane parameter $z_m = \sqrt{\lambda} / 2\pi m$; the size of gluonic cloud around this constituent quark is $z_m$ (see \cite{Chernicoff:2008sa} and references therein).

As $F \to F_{crit}$, the energy loss of the quark jet according to \eq{radpar} and \eq{critical} is given by 
 \beq
\frac{d E_{rad}}{d t} = F_{crit}
\eeq 
and is due to the string fragmentation, i.e. creation of quark--antiquark pairs from the vacuum. The force acting on the quark thus arises from the polarization of the vacuum by the "supercritical" charge which is screened by the creation of quark--anti-quark pairs from the vacuum -- this is the mechanism of quark confinement 
proposed by Gribov \cite{Gribov:1999ui}; for a review see \cite{Dokshitzer:2004ie}. We thus have arrived at the interpretation of confinement force at zero temperature as being due to the energy loss of supercritically charged quarks in the vacuum; this interpretation provides a simple relation between the mass of the produced constituent quark $m$, the coupling $\lambda = g^2 N_c = 12 \pi \as$ in $N_c = 3$ QCD, and the string tension $\sigma$:
\beq\label{tension}
F_{crit} =   \frac{2 \pi \ m^2}{\sqrt{\lambda}} = \sigma. 
\eeq
Using $m = 300$ MeV and $\as = 0.3$ (corresponding to $\sqrt{\lambda} \simeq 3.4$), we get for the string tension $\sigma \simeq 0.85$ GeV/fm -- quite reasonable value consistent both with phenomenology and the lattice data. The corresponding size of the constituent quark is $z_m = \sqrt{\lambda} / 2\pi m \simeq 0.35 \ {\rm fm}$. Of course, ${\cal N} = 4$ SUSY Yang-Mills theory is not confining; but once confinement is introduced as an external force, the formula \eq{tension} can tell us whether this force is "critical" and would result in the production of constituent quark pairs. The numerical estimates performed above suggest that the confinement force is indeed "critical".

\medskip

The case of quark acceleration parallel to the velocity $\vec a \parallel \vec v$ applies to the fragmentation of the jets in vacuum, and to the energy loss in the direction parallel to the colliding beams in hadron collisions. However it does not apply to jets produced around mid-rapidity in nuclear collisions. High energy collisions are accompanied by the creation of strings, or longitudinal color fields, which would exert a force perpendicular to the velocity of the jet produced at mid-rapidity. Such longitudinal fields of "supercritical" strength (i.e. capable of producing quark-antiquark and gluon pairs) have been shown to emerge in heavy ion collisions  \cite{Kharzeev:2005iz,Kharzeev:2006zm,Lappi:2006fp,Fries:2006pv,Iwazaki:2007es,Fujii:2008dd,Dumitru:2008wn} from the saturated parton distributions \cite{Gribov:1984tu,Blaizot:1987nc} in the color glass condensate \cite{McLerran:1993ni}. The produced longitudinal chromo-electric and chromo-magnetic fields (termed "glasma" in \cite{Lappi:2006fp}) possess non-zero Chern-Simons number  \cite{Kharzeev:2000ef,Shuryak:2001cp,Kharzeev:2001vs,Lappi:2006fp,Janik:2002nk}; the corresponding fluctuations of Chern-Simons number have been measured in real-time Yang-Mills calculation on the lattice \cite{Kharzeev:2001ev}.  The presence of Chern-Simons number can induce in heavy ion collisions the violation of parity \cite{Kharzeev:1998kz} (the possibility of spontaneous ${\cal T}$ and ${\cal P}$ violations has been considered in  \cite{Lee:1973iz}; in the context of heavy ion collisions, it has also been discussed by \cite{Morley:1983wr}). 
The ${\cal P}$ violation has been predicted to have a distinct signature \cite{Kharzeev:2004ey,Kharzeev:2007tn,Kharzeev:2007jp} -- charge asymmetry with respect to the reaction plane, resulting in the electric dipole moment of the produced quark-gluon matter; see \cite{Warringa:2008kv} for an overview. Recent preliminary experimental results indicate that this phenomenon may be present in RHIC data  \cite{Voloshin:2008kc}.
The propagation of charge in external classical fields of the type considered above has been considered by Shuryak and Zahed  \cite{Shuryak:2002ai}, who have evaluated synchrotron-like radiation basing on the extension of electrodynamics treatment due to Schwinger \cite{schwinger} to strong coupling. 

\medskip 

The longitudinal color fields will result in the acceleration perpendicular to the velocity, $\vec a \perp \vec v$. In this case the formulae \eq{lien} or \eq{relat} yield for the radiation power
\beq\label{radper}
\frac{d E_{rad}}{d t} = \frac{\sqrt{\lambda}}{2 \pi} \frac{1}{m^2}\ \gamma^2 \vec F^2.
\eeq 
This expression differs from \eq{radpar} in one but very important way -- it is proportional to the square of quark's energy $E = \gamma m$ (note that this will be true for any finite angle between $\vec a$ and $\vec v$).
Eq(\ref{radper}) is of course well known in classical electrodynamics \cite{Landau,Jackson} where it describes for example the energy loss due to synchrotron radiation. However in classical electrodynamics  \eq{radper} has to be supplemented by the condition on the strength of the field so that the classical description still applies \cite{Landau,Jackson}. Namely, the field strength $G_{rest}$ (we use this notation to avoid the confusion with the force $F$) in the rest frame of the charge  must obey the inequality 
\beq\label{cond}
G_{rest} \ll \frac{m^2}{e^3}
\eeq
In the frame where the charge moves with the velocity $v$, the field strength is $G = G_{rest}/\gamma$, and the Lorentz force acting on the charge is $F = e G$. Therefore the condition \eq{cond} translates into the following condition on external force $F$:
\beq\label{condforce}
F \ll \frac{m^2}{\gamma e^2}
\eeq
or in terms of external field $G$
\beq\label{condfield}
\gamma \frac{e^3 G}{m^2} \ll 1.
\eeq
As emphasized in \cite{Landau}, this condition does not prevent the ratio of the "radiation drag" force \eq{radper} to dominate over the external Lorentz force $F$; their ratio (we have substituted \eq{subst} for electrodynamics) is proportional to  
\beq
\frac{e^2}{m^2} \ \gamma^2\ F
\eeq
and at large energies $\gamma \gg 1$ will grow large even if \eq{condforce} is satisfied.  One is therefore justified to assume that the radiation drag force \eq{radper} is the dominant force acting on a relativistic particle at $\gamma \gg 1$. 
 
Shuryak and Zahed \cite{Shuryak:2002ai} have argued that in QCD the fields $G$ are strong, the coupling $e$ is large, and the external charge is massless, so the condition \eq{condfield} does not apply. They have thus concluded that Li\'enard formula \eq{lien} cannot be used in QCD \cite{Shuryak:2002ai}. However, the fact that equation \eq{lien} appears as the answer in the strong coupling relativistic problem in the AdS/CFT approach encourages us to take Li\'enard formula seriously. Indeed, the formula \eq{lien} holds in the ultra-relativistic limit; the coupling constant $\sqrt{\lambda}$ which replaces $e^2$ has been assumed large in the derivation \cite{Mikhailov:2003er}; the mass $m$ as discussed above has to be understood as a constituent mass; and the external force in AdS/CFT is limited only by the condition \eq{critical}. As emphasized above, Li\'enard formula relies only on locality, Lorentz invariance and dimensional analysis, and thus can be expected to have a wider range of validity than classical electrodynamics.

\medskip

Let us now come to the main point of this paper. As was noted by Pomeranchuk  \cite{chuk,Landau}, the formula \eq{radper} has a very interesting implication\footnote{Pomeranchuk considered the radiative energy loss of cosmic ray electrons in magnetic field of Earth \cite{chuk}.}. Indeed, the power of radiation is equal to the rate of energy loss, $dE_{rad}/dt = - dE/dt = -dE/dx$, where the last equality holds for a relativistic particle with $v \simeq c$. We thus can re-write \eq{radper} as 
\beq
- \frac{dE}{dx} = \frac{\sqrt{\lambda}}{2 \pi}\ \frac{{\vec F}^2(x)}{m^4}\ E^2,
\eeq
where we have explicitly indicated the dependence of the external force $F$ on the coordinate along the path. 
This differential equation can be easily solved by noting that $dE/E^2 = - d(1/E)$, and integrating over the path of length $L$ we get
\beq
\frac{1}{E_f} = \frac{1}{E_0} + \frac{\sqrt{\lambda}}{2 \pi}\ \int_0^L dx \ \frac{{\vec F}^2(x)}{m^4}.
\eeq
As the initial energy of the quark $E_0$ increases and goes to infinity, the final energy of the parton escaping from matter tends to a constant value! 

\medskip

Replacing $\int dx F^2(x)$ by the product of the path length $L$ and an average value $F^2$, we get for the upper bound on the energy of the parton 
\beq\label{bound}
E_{bound} = \frac{2 \pi}{\sqrt{\lambda}}\ \frac{m^4}{F^2}\ \frac{1}{L}.
\eeq
Let us measure the magnitude of the external force in units $z$ of the critical one given by \eq{critical}, $F \equiv z F_{critical}$.
We then get
\beq\label{simple}
E_{bound} = \frac{\sqrt{\lambda}}{2 \pi}\ \frac{1}{z^2 L}.
\eeq
If we assume as before in our discussion of the string tension that the external force is critical, $z \to 1$, we get from \eq{simple} a formula which depends only the value of the coupling and the path length:
\beq\label{simple1}
E_{bound} = \frac{\sqrt{\lambda_c}}{2 \pi}\ \frac{1}{L};
\eeq
where $\lambda_c$ is the critical value of the coupling corresponding to \eq{critical};
this formula applies when the masses of the propagating quark and the quarks produced from the vacuum are the same, i.e. when the propagating parton is light.

\medskip

In a more general case we can measure the magnitude of the force acting on a quark of mass $m$ in units $z_f$ of the critical value \eq{critical} for the creation of quarks of flavor $f$,
\beq
F = z^f\ F^f_{crit} = z^f\ \frac{2 \pi}{\sqrt{\lambda}}\ m_f^2; \ \ \ 0 \leq z^f \leq 1,
\eeq
the formula \eq{bound} then becomes
\beq\label{heavy}
E_{bound} = \frac{\sqrt{\lambda}}{2 \pi}\ \left(\frac{m}{m_f}\right)^4\ \frac{1}{z_f^2}\ \frac{1}{L}
\eeq

\medskip

Let us now make some numerical estimates. 
Assume first that the magnitude of the external force $F$ is given by the string tension, see \eq{tension}.  As before, we choose  
$\sqrt{\lambda} \simeq 3.4$ corresponding to $\as \simeq 0.3$, and let $z \to 1$; we then get from \eq{simple1}
\beq\label{estlight}
E_{bound}^{light} \simeq \frac{0.1}{L({\rm fm})}\ GeV;
\eeq
this means that none of the light high transverse momentum partons would escape from the longitudinal string\footnote{Note that we are discussing the case of strong coupling, and so perturbative processes are beyond the realm of this approach.}. For charm quarks, taking $m = 1.3$ GeV, $m_f = 0.3$ GeV and $z_f \to 1$ in  \eq{heavy}, we get
\beq
E_{bound}^{charm} \simeq \frac{35}{L({\rm fm})}\ GeV.
\eeq

Now let us take account of the fact that heavy ion collisions can produce much stronger color fields than encoded in \eq{tension}, as discussed above. It has been estimated that the magnitude of the produced fields at mid-rapidity is not much lower the critical one needed for the production 
of charm quarks \cite{Kharzeev:2003sk}. This leads us to take $m_f = m_c$, $z_c \simeq 1$ in \eq{heavy}; we then get for charm the same estimate as we got above for light partons, \eq{estlight}
\beq
\tilde{E}_{bound}^{charm} \simeq \frac{0.1}{L({\rm fm})}\ GeV. 
\eeq
and for beauty with $m_b \simeq 4.5$ GeV
\beq\label{bquark}
\tilde{E}_{bound}^{beauty} \simeq \frac{14}{L({\rm fm})}\ GeV. 
\eeq
We should of course admit that the estimates above are very rough, and depend 
crucially on the magnitude of the produced color fields.
\medskip

These estimates suggest that in strong color fields produced in relativistic heavy ion collisions at RHIC both light and charm quarks 
and gluons cannot escape from dense region of the produced matter. 
Of course this does not mean that there will be no high transverse momentum particles -- they will be emitted from the surface of the produced fireball, leading to a universal normalized ratio of nuclear and proton-proton cross sections $R_{AA}$ for gluon, light and charm quarks almost independent of energy. 
A weak increase of $R_{AA}$ may result from the small amount of conventional absorption in the dilute "corona" surrounding the dense core of the plasma which is expected to go away at large transverse momentum in accord with factorization theorems of perturbative QCD. This leaves little room for observing the medium-induced modifications of the jet structure -- either the jet is produced in the "corona" and is thus not modified at all, or it is produced in the dense core and is completely absorbed, with the final energy
below the bounds given above. 

Based on the perturbative arguments, 
one is led to search for the jet modification in the central collisions of heavy nuclei. The bound presented here is inversely proportional to the length of traversed medium, and so suggests that the only hope to observe 
the jet modified by the medium-induced radiation is in peripheral 
collisions and/or in the collisions of lighter ions. The color fields, and thus the external force acting on the color charge, are proportional to the saturation momentum squared, $F \sim Q_s^2$ which grows with the centrality as $Q_s^2 \sim N_{part}^{1/3}$ where $N_{part}$ is the number of participant nucleons (see \cite{Kharzeev:2000ph} for details). Since $L \sim N_{part}^{1/3}$ the bound may be 
expected to depend on centrality as $E_{bound} \sim 1/N_{part}$.
Of course, detailed calculations including the realistic nuclear geometry and taking account of the trigger bias effect enhancing the 
contribution of the smaller path lengths will have to be performed to get a reliable estimate of the bound.

The color fields, and thus the external force acting on the color charge, are proportional to the saturation momentum squared $F \sim Q_s^2$ which grows with energy the bounds will decrease at the LHC energy. According to the estimates of the saturation momentum at the LHC and RHIC (see e.g. \cite{Kharzeev:2004if}), the bound on the $b$ quarks at the LHC will therefore change to
\beq\label{bquark}
\tilde{E}_{bound}^{beauty} \simeq \frac{2}{L({\rm fm})}\ GeV, 
\eeq 
so the suppression of $b$ quarks at high $p_{\perp}$ should become a clearly visible effect.

\medskip

Let us note also that once the massive constituent partons are slowed down below the energy given by the bound, their velocity will become significantly smaller than the velocity of light. Therefore the emitted radiation must possess a "dead cone"of size $\theta_{cone} \simeq m/E$. The energy $E$  of the radiating parton will decrease from the initial energy to the final energy given by the bound. To get an upper bound on the size of the dead cone we use $E = E_{bound}$ and get
\beq\label{cone}
\theta_{cone} < \frac{m}{E_{bound}} \simeq \frac{2 \pi}{\sqrt{\lambda}}\ m L,
\eeq
where $m$ is the constituent mass; we have used the expression \eq{simple1} for the bound on the light parton energy. Since the emitted energy-momentum is linear in the path of the parton, the resulting angular distribution will be a linear superposition of emissions from partons with energies ranging from some initial energy $E_0$ to the final energy $E_{bound}$, with "dead cones" ranging from 
very small $\theta_{cone} \simeq m/E_0$ to the large value \eq{cone}, $m/E_0 < \theta_{cone} < m/E_{bound}$.
Therefore the resulting dead cone can be quite broad; again, detailed calculations are needed to evaluate the shape of the angular distribution of the emitted radiation.

\medskip

To summarize, we have shown that if Li\'enard formula derived through the AdS/CFT correspondence holds in the strong coupling regime in QCD then 
under quite natural assumptions about the evolution of gauge fields in heavy ion collisions there should exist an upper bound on the final energy of the parton escaping from the strongly coupled matter.  
Of course, the asymptotic freedom of QCD dictates that at some large transverse momentum the dynamics should become perturbative. However it is not yet clear 
at what scale the strong and weak coupling descriptions match in heavy ion collisions; an experimental study of the possible existence of the bound  is needed to answer this question. 

\medskip

 This work was supported by the U.S. Department of Energy under Contract No. DE-AC02-98CH10886.


\end{document}